Solve the linear inhomogeneous differential equation using the constant variation method: y,-2y/x=I +1/x


1
Expert's answer
2021-11-30T07:25:32-0500

Corresponding homogeneous differential equation


"y'-2y\/x=0"

"\dfrac{dy}{y}=2\dfrac{dx}{x}"

Integrate


"\int\dfrac{dy}{y}=\int2\dfrac{dx}{x}"

"y=Cx^2"

"y'=C'x^2+2xC"

"C'x^2+2xC-2\dfrac{Cx^2}{x}=1+\dfrac{1}{x}"

"C'=\dfrac{1}{x^2}+\dfrac{1}{x^3}"

Integrate


"C=-\dfrac{1}{x}-\dfrac{1}{2x^2}+c_1"

Substitute


"y=-x-\dfrac{1}{2}+c_1x^2"